OpenStudy (anonymous):

Multiple choice: Can someone please please help me with some more surface area of pyramids and cones???

OpenStudy (anonymous):

a) 20 b) 10.6 C08.9 d) 14.4 e)7.1

OpenStudy (anonymous):

20

OpenStudy (anonymous):

A=1/2(LR)+pi*r^2 ==>138.23=1/2(2*pi*2)*R+pi*2^2 R=19.999

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

find the surface area of the solid. a)716.3m^2 b)867.1m^2 c)1168.7m^2 d)1055.6m^2 e)942.5m^2